OpenStudy (anonymous):

A school fundraiser sells holiday cars and wrapping paper. They are trying to raise at least $400. They make a profit of $1.50 on each box of holiday cards and $1.00 each pack of wrapping paper. A. What is an inequality for the profit the school wants to make for the fundraiser? B.If the fundraiser sells 100 boxes of cars and 160 packs of wrapping paper, will they reach their goal? Show your work.

OpenStudy (anonymous):

h = holiday cards w = wrapping paper A.) 1.5h + 1w >= 400 B.) 1.5(100) + 1(160) <= 400 150 + 160 >= 400 310 > = 400 (false) The school fundraiser, under these conditions, will not reach it's goal.